$SPAD/src/input fnla.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fnla.output
)set message test on
)set message auto off
)clear all
--S 1 of 7
fnl := FNLA(4,4,INT)
--R
--R
--R   (1)  FreeNilpotentLie(4,4,Integer)
--R                                                                 Type: Domain
--E 1
--S 2 of 7
dimension()$fnl
--R
--R
--R   (2)  90
--R                                                     Type: NonNegativeInteger
--E 2
--S 3 of 7
x:fnl := generator(1)
--R
--R
--R   (3)  e
--R         1
--R                                          Type: FreeNilpotentLie(4,4,Integer)
--E 3
--S 4 of 7
y:fnl := generator(17)
--R
--R
--R   (4)  e
--R         17
--R                                          Type: FreeNilpotentLie(4,4,Integer)
--E 4

--S 5 of 7
z:=x*y
--R
--R
--R   (5)  2e   - e   + e
--R          78    45    38
--R                                          Type: FreeNilpotentLie(4,4,Integer)
--E 5
--S 6 of 7
deepExpand z
--R
--R
--R   (6)  2[[e ,e ],[e ,e ]] - [e ,[e ,[e ,e ]]] + [e ,[e ,[e ,e ]]]
--R            1  2    2  3       3   2   1  2        2   2   1  3
--R                                                             Type: OutputForm
--E 6
--S 7 of 7
shallowExpand z
--R
--R
--R   (7)  2[e ,e ] - [e ,e  ] + [e ,e  ]
--R           5  8      3  14      2  15
--R                                                             Type: OutputForm
--E 7
)spool
)lisp (bye)

———-

References^{[1] nothing}

5